OpenStudy (anonymous):

A hyperbolic mirror can be used to take panoramic photos, if the camera is pointed toward the mirror with the lens at one focus of the hyperbola. Write the ewuation of the hyperbola that can be used to model a mirror that has a vertex 4 inches from the center of the hyperbola and a focus 1 inch in front of the surface of the mirror. Assume the mirror has a horizontal transverse axis and the hyperbola is conetered at (0,0). Is the vertes (+-4,0) and the focus (5,0)?

OpenStudy (dumbcow):

if i am reading it correctly, then yes you are correct vertices (+-4,0) foci (+-5,0)

OpenStudy (anonymous):

I was doing foci (1,0) but decided I needed to add the 1 to the 4 to get (5,0).a^2=16, b^2=9 and c^2=1 right?

OpenStudy (dumbcow):

no c^2 = 25 because "c" is distance focus is from center

OpenStudy (dumbcow):

but you were right to change it :)

OpenStudy (anonymous):

right thanks

OpenStudy (anonymous):

x^2/16-y^2/9=1